OpenStudy (anonymous):

Solve the following equation. Please show all of your work. 2 - 3/4*x=20

sam (.sam.):

((2-3)/4 )*x=20 or 2-3x/4=20 ?????????

OpenStudy (anonymous):

the first

sam (.sam.):

((2-3)/(4))*x=20 ((-1)/(4))*x=20 (-(1)/(4))*x=20 x(-(1)/(4))=20 -(x)/(4)=20 -x=80 x=80*-1 x=-80